6)

Two charges of q_1 = 1.2 muC and q_2 = -2.0 are d
Answer #1

(a) The electric potential due to the 1.2 \muC charge at the third vertex, point P will be given as :

V1 = ke q1 / d = (9 x 109 Nm2/C2) (1.2 x 10-6 C) / (0.66 m)

V1 = 1.63 x 104 V

(b) The electric potential due to the -2 \muC charge at the third vertex, point P will be given as :

V2 = ke q2 / d = (9 x 109 Nm2/C2) (2 x 10-6 C) / (0.66 m)

V2 = -2.72 x 104 V

(c) The total electric potential at P which will be given as :

\DeltaV = V1 + V2 = [(1.63 x 104 V) - (2.72 x 104 V)]

\DeltaV = - 1.09 x 104 V

(d) The work required to move a 3.5 \muC charge from an infinity to P which will be given as :

using a formula, we have

\DeltaW = q \DeltaV = (3.5 x 10-6 C) (-1.09 x 104 V)

\DeltaW = - 3.81 x 10-2 J